A 66 kg gymnast is at the top of a swing on the high bar. Her velocity is 2.3 m/s and she is 3.8 m high. What is the gymnast's total mechanical energy of this instant?
Answer #1

The kinetic energy is given by KE = 1/2 mv^2 = 174.57 J

The potential energy is given by PE = mgh = 2457.84 J

The total mechanical energy is given by E = KE + PE = 2632.41 J